# Efficient optimization of neural network backflow for ab-initio quantum chemistry

### v1

@@ II.4 Gumbel top-k trick
A natural candidate for applying the Gumbel top-$k$ trick is the preselected subspace $\tilde{\mathcal{U}}$ from Section II.2. Unlike the original FSSC scheme, which relies on the dominance of the core space $\mathcal{V}$, using the Gumbel top-$k$ trick allows us to unbiasedly represent $\tilde{\mathcal{U}}$, which is more representative than $\mathcal{V}$ alone, as it is preselected for importance and grows extensively with system size when |$\absolutevalue{\mathcal{V}}$ | is fixed. By introducing Gumbel noise, previously excluded borderline configurations can be sampled, improving the approximation of both energy and gradients. Moreover, the computational overhead is minimal since Gumbel noise generation is inexpensive.
To formalize this approach, we describe how the Gumbel top-$k$ trick is used to form the sample $\tilde{\mathcal{S}}$ and assign importance weights $\tilde{w}_{i}$ for constructing more unbiased estimates of equation (1) and (3). The new core space can still be deterministically selected as in Ref. 34 or simply chosen as $\tilde{\mathcal{S}}$. Samples are drawn from $\tilde{\mathcal{U}}$ by perturbing the log probability [Eq. (6)] with independently and identically distributed Gumbel noise:
@@
and selecting the top-$k$ configurations with perturbed log-probabilities:
$\tilde{\mathcal{S}}=\underset{\ket{\mathbf{x}_{i}}\in\tilde{\mathcal{U}}}{\text{argtop}k},[G_{i}=\log\tilde{p}_{i}+g_{i}].$ (10)
The corresponding importance weight is then given as:
$\tilde{w}_{i}(\kappa)=\frac{\tilde{p}_{i}}{q_{i}(\kappa)}=\frac{\tilde{p}_{i}}{1-e^{-e^{\log\tilde{p}_{i}-\kappa}}},$ (11)
where $\kappa$ is the $(k+1)$-th largest $G_{i}$, serving as the empirical threshold (since |$\ket{\mathbf{x}_{i}}\in\tilde{\mathcal{S}}$ if $G_{i}>\kappa$), and $q_{i}(\kappa)=P(G_{i}>\kappa)$ denotes the inclusion probability of selecting sample $i$. Importance weights are typically renormalized to reduce variance in practice [42], albeit at the cost of introducing bias.
Using these weights and the Gumbel-top-$k$-selected samples $\tilde{\mathcal{S}}$, we construct improved estimates for the energy and gradient:
$E_{\theta}\approx\sum_{\ket{\mathbf{x}_{i}}\in\tilde{\mathcal{S}}}\tilde{w}_{i}(\kappa)\tilde{E}_{l}(\mathbf{x}_{i}|\tilde{\mathcal{U}})$ (12)
@@
where $\tilde{E}_{l}(\mathbf{x}_{i}|\tilde{\mathcal{U}})$ is equation (8). While equation (12) and (13) are not fully unbiased estimators of equation (1) and (3), they significantly reduce bias compared to the deterministic FSSC approach.
Figure 2: Comparison of the energy optimization curve and the number of unique configurations sampled per iteration between the FSSC ($N_{s}=1024$) and MCMC ($N_{s}=32768$) schemes for Li2O, using the STO-3G basis set and canonical HF orbitals. Solid lines represent the objective energy optimization curve, while dashed lines indicate the number of unique configurations sampled per iteration. The blue star marks the post-training MCMC inference energy for the FSSC scheme. A moving average window of 400 is applied for improved readability.
@@ II.5.1 Enforce Spin-flip symmetry
This construction ensures both $\hat{F}\ket{\psi_{SFS,\theta}}=\pm\ket{\psi_{SFS,\theta}}$ and $\psi_{SFS,\theta}(\hat{F}\mathbf{x})=\pm\psi_{SFS,\theta}(\mathbf{x})$. Unlike the approach in Ref. [28], which enforces a weaker form of the symmetry (|$\absolutevalue{\psi(\hat{F}\mathbf{x}_{i})}=\absolutevalue{\psi(\mathbf{x}_{i})}$ |) through specialized preprocessing of subnetwork inputs and postprocessing of outputs, our method is more general. It requires only the addition of a single equation atop any wavefunction ansatz, ensuring full spin-flip symmetry while naturally preserving the relative sign consistency between spin-flip-equivalent configurations.
@@ II.5.2 Spin-flip-symmetry-aware strategies
When a wavefunction ansatz is constructed as in Section II.5.1, several strategies can be leveraged to enhance the training pipeline.
For constructing the preselected subspace $\tilde{\mathcal{U}}$, among each pair of spin-flip-equivalent configurations (or the single representative in the spin-flip-symmetric case), only the one with the higher importance score should be retained, as both share identical amplitude magnitudes.
During sample generation via the Gumbel top-$k$ trick, non-spin-flip-symmetric configurations |$\ket{\mathbf{x}_{i}}$ have the same probability and local observables (e.g., local energy) as their excluded counterparts $\hat{F}\ket{\mathbf{x}_{i}}$. This redundancy can be utilized by doubling the probability of non-symmetric configurations:
$\tilde{p}_{\text{SFS},i}=\begin{cases}2\tilde{p}_{i},&\text{if }\ket{\mathbf{x}_{i}}\text{ is non-spin-flip-symmetric},\\ \tilde{p}_{i},&\text{if }\ket{\mathbf{x}_{i}}\text{ is spin-flip-symmetric}.\end{cases}$ (15)
Similarly, when evaluating the truncated local energy [Eq. (8)], the lookup for the precomputed amplitude $\psi_{\theta}(\mathbf{x}_{j})$ in $\tilde{\mathcal{U}}$ should also include $\psi_{\theta}(\hat{F}\mathbf{x}_{j})$.
These modifications effectively double the sample size without compromising the efficiency of the truncated local energy strategy, as spin-flip-symmetric configurations are significantly fewer compared to their non-symmetric counterparts.
@@ II.5.3 Orbital occupation
To incorporate this prior knowledge, we define an ordered position representation of the configuration bitstring as $\mathbf{y}(\mathbf{x})=vector(\{i|x^{i}=1\})$, which lists the indices of occupied orbitals in ascending order. Using this representation, we propose the following trainable discrete orbital envelope:
$\pi_{\mathbf{\alpha}}(\mathbf{y}(\mathbf{x}))=$,
where $\alpha$ is a set of trainable parameters. Each term acts as a penalty based on the displacement of the occupied orbital from its reference position, with lower-indexed orbitals favored for lower-indexed electrons. To respect spin-flip symmetry, both spin channels share the same set of envelope parameters $\alpha$. See Appendix B for a visualization and analysis of the learned orbital envelope parameters after training.
@@ III Results
We evaluate the performance of NNBF combined with the algorithmic enhancements proposed in Section II.2, II.3, II.4, and II.5 on various molecules using the STO-3G basis set as well as on a paradigmatic strongly correlated system: frozen-core N2 with the cc-pVDZ basis set. To demonstrate the improvements from these enhancements, we compare results with our previous algorithm [34] and conduct a thorough ablation study by cumulatively adding the proposed techniques. We also investigate the relationship between NNBF’s representability and the inverse participation ratio (IPR). Details on the default network architectures, hyperparameters, training routine, and post-training MCMC inference procedure are provided in Appendix A.
@@ III.1 Benchmarks
We assess the performance of our algorithms by first comparing molecular ground-state energies with CCSD and CCSD(T) baselines and other NQS results, using molecular geometries sourced from PubChem [44]. All calculations follow the protocols described in Sections II.2, II.3, II.4, and II.5, with a fixed batch size $N_{s}=1024$, which is equal to the core space size (|$\absolutevalue{\mathcal{V}}$ |). Table 1 shows that NNBF, using only default hyperparameters and the improved algorithms, not only outperforms conventional CCSD methods but also achieves lower energies than all existing NQS approaches, particularly excelling in larger molecular systems. For smaller molecules where FCI energies exist, it’s often the case that NNBF is closer than CCSD(T) to the exact answer. In our largest two molecules, there is no FCI energy to compare against. In these cases, we find the energy is very close to CCSD(T); because CCSD(T) isn’t variational, it is unclear in those cases whether NNBF or CCSD(T) is better.
To further benchmark our method, we calculate the ground-state energy of frozen-core N2 using the cc-pVDZ basis. As shown in Table 2, with canonical HF orbitals, NNBF achieves superior variational energies compared to SCHI [10] and ASCI [12], which represent the state of the art in SCI methods for QC. After orthonormally rotating the unfrozen orbitals—an operation that does not affect the ground-state calculation for either frozen-core or all-electron treatments—NNBF attains variational energies that are lower but still within error bars (approximately 0.1 mHa difference) of the best results from SCHI, ASCI, and FCIQMC [38], underscoring its ability to capture strongly correlated systems. Other NQS methods [35] yield energy differences on the order of several mHa, further highlighting the superiority of the NNBF ansatz. These results are obtained using a moderate batch size ($N_{s}=8192$) and a neural network with $h=512$ hidden units.
To demonstrate the impact of our algorithmic improvements, we compare the current algorithm with that of Ref. 34 by studying how the batch size used during optimization affects NNBF performance on Li2O with the STO-3G basis set and canonical HF orbitals. Figure 3 shows that while both methods exhibit polynomial scaling of energy improvement with batch size, our updated algorithm achieves a faster convergence rate of $O(N_{s}^{-1.979616\pm 0.006350})$ compared to $O(N_{s}^{-1.166309\pm 0.014370})$ for the previous method. Notably, our new approach requires orders of magnitude fewer samples to reach the same energy accuracy. Furthermore, the wall time per optimization step is significantly reduced—details are discussed in Section III.2.
Molecule $N_{t}$ CCSD CCSD(T) FCI Best NQS NNBF
@@
CH4O $4.01\times 10^{6}$113.665485 -113.666214 -113.666485 -113.665485[29] -113.666379(76)
Li2O $4.14\times 10^{7}$87.885514 -87.893089 -87.892693 -87.8922[31] -87.892689(25)
C2H4O $2.54\times 10^{9}$151.120474 -151.122748 - -151.12153[31] -151.122602(39)
C2H4O2 $5.41\times 10^{11}$225.050896 -225.057823 - -225.0429767[29] -225.055531(75)
@@
Table 1: Benchmarking NNBF ground-state energies against conventional CCSD, CCSD(T), and FCI results, as well as the best published NQS results, with footnotes indicating the respective methods (excluding our previous work [34]). $N_{t}$ represents the total number of physically valid configurations that conserve both particle number and total charge. For NNBF, we run three different training runs each following the default settings specified in Appendix A with $N_{s}=1024$, and the NNBF energy column reports the best post-training MCMC inference energy.
Method Number of determinants Variational energy Total energy
HF orbitals
SHCI[10] 37593 -109.2692 -109.2769
ASCI[12] 10000 -109.26419 -109.27687
ASCI[12] 30000 -109.26936 -109.27691
ASCI[12] 100000 -109.27335 -109.27698
FCIQMC[38] - - -109.2767(1)
NNBF 8192 -109.27616(5) -
MP2 natural orbitals
NNBF 8192 -109.2771(1) -
Dynamic orbital rotations
ASCI[12] 10000 -109.26837 -109.27708
ASCI[12] 100000 -109.27522 -109.27699
ASCI[12] 300000 -109.27638 -109.27699

### v2

@@ II.4 Gumbel top-k trick
A natural candidate for applying the Gumbel top-$k$ trick is the target space $\mathcal{U}$ from Section II.2. Unlike the original FSSC scheme, which relies on the dominance of the core space $\mathcal{V}$, using the Gumbel top-$k$ trick allows us to unbiasedly represent $\mathcal{U}$, which is more representative than $\mathcal{V}$ alone, as it is constructed for importance and grows extensively with system size when |$\absolutevalue{\mathcal{V}}$ | is fixed. By introducing Gumbel noise, previously excluded borderline configurations can be sampled, improving the approximation of both energy and gradients. Moreover, the computational overhead is minimal since Gumbel noise generation is inexpensive.
To formalize this approach, we describe how the Gumbel top-$k$ trick is used to form the sample $\mathcal{S}$ and assign importance weights $w_{i}$ for constructing more unbiased estimates of equation (1) and (3). The procedure begins with a normalized probability distribution over the target space $\mathcal{U}$, derived from the precomputed amplitudes $\psi_{\theta}(\mathbf{x}_{i})\ \forall\ket{\mathbf{x}_{i}}\in\mathcal{U}$:
$p_{i}=\frac{\psi^{2}_{\theta}(\mathbf{x}_{i})}{\sum_{\ket{\mathbf{x}_{j}}\in\mathcal{U}}\psi^{2}_{\theta}(\mathbf{x}_{j})},\quad\forall\ket{\mathbf{x}_{i}}\in\mathcal{U}.$ (7)
Samples are drawn from $\mathcal{U}$ by first perturbing this log probability with independently and identically distributed Gumbel noise:
@@
and then selecting the top-$k$ configurations with perturbed log-probabilities:
$\mathcal{S}=\underset{\ket{\mathbf{x}_{i}}\in\mathcal{U}}{\text{argtop}k}[G_{i}=\log{p}_{i}+g_{i}].$ (9)
To construct an unbiased estimator over the target space $\mathcal{U}$, we follow the Horvitz-Thompson estimator by assigning each sample an importance weight, $w_{i}$, defined as the ratio of its original probability $p_{i}$ to its inclusion probability $q_{i}(\kappa)$[43, 42]. The inclusion probability $q_{i}(\kappa)$—the probability of configuration $i$ being selected—is determined by an empirical threshold $\kappa$ set by the $(k+1)$-th largest perturbed log-probability, $G_{i}$. The formulas for the inclusion probability and the final weight are:
$q_{i}(\kappa)$ $=P(G_{i}>\kappa)=1-e^{-e^{(\log{p}_{i}-\kappa)}}$ (10)
$w_{i}(\kappa)$ $=\frac{p_{i}}{q_{i}(\kappa)}=\frac{p_{i}}{1-e^{-e^{(\log{p}_{i}-\kappa)}}}$ (11)
Importance weights are typically renormalized over sample $\mathcal{S}$, i.e. $w_{i}\leftarrow w_{i}/\sum_{j\in\mathcal{S}}w_{j}$, to reduce variance in practice [43], albeit at the cost of introducing bias. The effects of the Gumbel noise, the use of inclusion probabilities for weighting, and weight renormalization are investigated in detail in Section III.2.2.
Using these weights and the Gumbel-top-$k$-selected samples $\mathcal{S}$, we construct improved estimates for the energy and gradient:
$E_{\theta}\approx\sum_{\ket{\mathbf{x}_{i}}\in\mathcal{S}}w_{i}(\kappa)E_{l}(\mathbf{x}_{i}|\mathcal{U})$ (12)
@@
where $E_{l}(\mathbf{x}_{i}|\mathcal{U})$ is equation (6). While equation (12) and (13) are not fully unbiased estimators of equation (1) and (3), they significantly reduce bias compared to the deterministic FSSC approach.
Figure 2: Comparison of the energy optimization curve and the number of unique configurations sampled per iteration between the FSSC [34] (|$\absolutevalue{\mathcal{S}}=1024$) and MCMC (|$\absolutevalue{\mathcal{S}}=32768$) schemes for Li2O, using the STO-3G basis set and canonical HF orbitals, where |$\absolutevalue{\mathcal{S}}$ | denotes the size of sample set. Solid lines represent the objective energy optimization curve, while dashed lines indicate the number of unique configurations sampled per iteration. The blue star marks the post-training MCMC inference energy for the FSSC scheme. A moving average window of 400 is applied for improved readability.
@@ II.5.1 Enforce Spin-flip symmetry
This construction ensures both $\hat{F}\ket{\psi_{SFS,\theta}}=\pm\ket{\psi_{SFS,\theta}}$ and $\psi_{SFS,\theta}(\hat{F}\mathbf{x}_{i})=\pm\psi_{SFS,\theta}(\mathbf{x}_{i})$. Unlike the approach in Ref. [28], which enforces a weaker form of the symmetry (|$\absolutevalue{\psi(\hat{F}\mathbf{x}_{i})}=\absolutevalue{\psi(\mathbf{x}_{i})}$ |) through specialized preprocessing of subnetwork inputs and postprocessing of outputs, our method is more general. It requires only the addition of a single equation atop any wavefunction ansatz, ensuring full spin-flip symmetry while naturally preserving the relative sign consistency between spin-flip-equivalent configurations.
@@ II.5.2 Spin-flip-symmetry-aware strategies
When a wavefunction ansatz is constructed with spin-flip symmetry, as described in Section II.5.1, several strategies can be leveraged to enhance the training pipeline.
The process of updating the subspace $\mathcal{U}$ (see Section II.2) can explicitly incorporate spin-flip symmetry. For any pair of spin-flip-equivalent configurations (or for a single configuration if it is self-spin-flip-equivalent), only one unique representative is considered to be selected into $\mathcal{U}$, determined by a predefined rule. This choice is justified because both configurations in such a pair share an identical amplitude magnitude. This initial filtering of $\mathcal{U}$ to remove spin-based redundancy provides several downstream advantages.
First, this efficiency propagates to the construction of the core and connected spaces. The two sets of configurations connected to any spin-flip-equivalent pair are themselves spin-flip equivalent. Therefore, if the core space $\mathcal{V}$ is selected from the now redundancy-free subspace $\mathcal{U}$, then the expansion of connected configurations from this new $\mathcal{V}$ will also avoid such spin-flip redundancies.
Second, during sample generation from $\mathcal{U}$ via the Gumbel top-$k$ trick, the treatment of non-spin-flip-symmetric configurations can be adjusted. Consider a configuration |$\ket{\mathbf{x}_{i}}\in\mathcal{U}$ that is part of a non-symmetric pair. Its partner, $\hat{F}\ket{\mathbf{x}_{i}}$, is not stored in $\mathcal{U}$. However, both configurations have the same probability ($p(\mathbf{x}_{i})=p(\hat{F}\mathbf{x}_{i})$) and the same local energy. To account for the excluded partner, we effectively transfer its importance weight to the representative configuration stored in $\mathcal{U}$ by doubling its sampling probability:
$p_{\text{SFS},i}=\begin{cases}2p_{i},&\text{if }\ket{\mathbf{x}_{i}}\in\mathcal{U}\text{ is non-spin-flip-symmetric},\\ p_{i},&\text{if }\ket{\mathbf{x}_{i}}\in\mathcal{U}\text{ is spin-flip-symmetric},\end{cases}$ (15)
where $p_{i}$ refers to the probabilities derived from Eq. (7). These modified probabilities are then used in the Gumbel top-$k$ sampling and can be renormalized as needed.
Third, this symmetry awareness extends to the evaluation of the truncated local energy (Eq. (6)). Although $\mathcal{U}$ stores amplitudes $\psi_{\theta}(\mathbf{x}_{i})$ only for unique representatives, the amplitude of a spin-flipped partner, $\psi_{\theta}(\hat{F}\mathbf{x}_{i})$, is also implicitly known through the symmetry relation $\psi_{\theta}(\hat{F}\mathbf{x}_{i})=\pm\psi_{\theta}(\mathbf{x}_{i})$. Thus, when computing local energies, all connected configurations are first transformed by the predefined rule. The amplitude information of these transformed connected configurations is then retrieved from $\mathcal{U}$, and the spin-flip-symmetry phase will be applied to the retrieved amplitude if the connected configuration was indeed spin-flipped. This effectively allows the local energy calculation to leverage amplitude information from an expanded set of configurations.
These modifications effectively increase the utilized sample size and information content—often doubling it, since true spin-flip-symmetric configurations are typically far less numerous than their non-symmetric counterparts—all without compromising the efficiency of the truncated local energy strategy.
@@ II.5.3 Orbital occupation
To incorporate this prior knowledge, we first define an ordered occupied-position representation of the configuration bitstring as $\mathbf{y}^{\uparrow/\downarrow}(\mathbf{x})=vector(\{i|x^{i\uparrow/\downarrow}=1\})$, which lists the indices of occupied spin up/down orbitals in ascending order. Using this representation, we propose the following trainable discrete orbital envelope:
$\pi_{\bm{\alpha}}(\mathbf{x})=\exp\Biggl(-\sum_{i=1}^{N_{e}/2}\left[|\alpha_{i}(y^{\uparrow}_{i}-i)|+|\alpha_{i}(y^{\downarrow}_{i}-i)|\right]\Biggr{missing}),$ (16)
where $\bm{\alpha}=\{\alpha_{i}\}$ is a set of trainable parameters. The term $|y^{\sigma}_{i}-i|$ measures the displacement of the $i$-th electron of spin $\sigma\in\{\uparrow,\downarrow\}$ from its reference orbital index $i$. Each parameter $\alpha_{i}$ learns a penalty for this displacement, effectively suppressing configurations where electrons are excited into higher-energy orbitals and favoring lower-indexed electrons for lower-indexed orbitals. To respect spin-flip symmetry, both spin channels share the same set of envelope parameters $\bm{\alpha}$. A visualization and analysis of the learned orbital envelope parameters after training are provided in Appendix B.
@@ III Results
We first evaluate the performance of NNBF combined with the algorithmic enhancements proposed in Section II.2, II.3, II.4, and II.5 on various molecules using the STO-3G basis set as well as on paradigmatic strongly correlated systems: the dissociation curve of H2O utilizing the cc-pVDZ basis set, and frozen-core and all-electron N2 molecule calculations also with the cc-pVDZ basis set at representative bond lengths.
To demonstrate the improvements from these enhancements, we present three distinct analyses. First,we conduct a thorough ablation study by cumulatively adding the proposed techniques starting from our previous algorithm [34]. Second, we perform a focused study on the Gumbel top-$k$ selection scheme described in Section II.4, examining the specific roles of the Gumbel noise, the use of inclusion probabilities, and weight renormalization. Finally, we directly compare our proposed truncated local energy strategy (Section II.3) against an approach commonly employed in the community [32, 35, 36].
The relationship between the representational capacity of NNBF and the inverse participation ratio (IPR) of the quantum state is also investigated. Specific details regarding the (default) neural network architectures, hyperparameters, training protocols, and the post-training MCMC inference procedure are provided in Appendix A.
@@ III.1 Benchmarks
III.1.1 Ground state energy for various molecules
We assess the performance of our enhanced NNBF algorithms by first comparing calculated molecular ground-state energies against those from established CCSD and CCSD(T) baselines, as well as results from other NQS methods. These calculations utilize molecular geometries sourced from PubChem [45], which is also provided in Table 5 in Appendix A.2 for reference and convenience, and strictly adhere to the computational protocols described in Section II.2, Section II.3, Section II.4, and Section II.5. The results, summarized in Table 1, demonstrate that NNBF employed with the improved algorithms not only generally outperforms conventional CCSD methods and achieves energies comparable or superior to CCSD(T) for many systems, but also consistently yields lower energies than other existing NQS approaches, particularly excelling for larger molecular systems.
Molecule $||\absolutevalue{\mathcal{H}}$ CCSD CCSD(T) FCI Best NQS NNBF
@@ III.1.1 Ground state energy for various molecules
CH4O $4.01\times 10^{6}$113.665485 -113.666214 -113.666485 -113.665485[29] -113.666416(26)
Li2O $4.14\times 10^{7}$87.885514 -87.893089 -87.892693 -87.8922[31] -87.892662(18)
C2H4O $2.54\times 10^{9}$151.120474 -151.122748 -151.123570 -151.12153[31] -151.123357(28)
C2H4O2 $5.41\times 10^{11}$225.050896 -225.057823 - -225.0429767[29] -225.058589(40)
@@
Table 1: A benchmark of ground-state energies from our NNBF method against conventional quantum chemistry methods (CCSD, CCSD(T), FCI) as well as the best published NQS results, with footnotes indicating the respective methods (excluding our previous work [34]). $|\mathcal{H}|$ is the size of the total Hilbert space, comprising all configurations that conserve both particle number and total spin projection ($S_{z}$). The reported NNBF energy is obtained via the following protocol: five independent training runs were performed (settings in Appendix A), and a post-training MCMC inference was used to estimate the energy of each. The single model with the lowest of these five energies was selected. A final, separate MCMC inference was then conducted on this best model to obtain the unbiased estimate reported in the table.
III.1.2 All electron H2O dissociation curve
To investigate the ability of the NNBF method to describe strong quantum correlations, we computed the dissociation curve for an all-electron H2O molecule using the cc-pVDZ basis set, with the bond angle held fixed at $104.5^{\circ}$. As illustrated in Figure 3, the resulting NNBF energies are in excellent agreement with the FCI benchmark, achieving a mean absolute error of only 0.08 mHa across the entire curve. Notably, NNBF outperforms conventional quantum chemistry approaches for both near-equilibrium geometries and stretched bond lengths. The latter is a region where the gold-standard CCSD(T) method is known to falter due to the increasing importance of static correlation at large bond separations. This result demonstrates the proficiency of NNBF in accurately capturing both static and dynamic electron correlations.
Figure 3: Dissociation curve of the H2O molecule, calculated using a cc-pVDZ basis set with the bond angle fixed at $104.5^{\circ}$. The plot compares the performance of our NNBF method against standard quantum chemistry approaches (HF, CISD, CCSD, CCSD(T)) and the exact FCI energy, which serves as the ground-truth reference. The NNBF results were obtained from a single training run per bond length with parameters $|\mathcal{S}|=8192$, $|\mathcal{V}|=2048$, and a $(D,L,h)=(1,2,512)$. Each reported NNBF energy is the result of a single post-training MCMC inference.
III.1.3 All-electron and frozen-core N2 calculations
To further evaluate our method against other state-of-the-art ab-initio techniques, we calculated the ground-state energy of frozen-core N2 with the cc-pVDZ basis set. The results, presented in Table 2, show that when using canonical Hartree-Fock orbitals, NNBF achieves superior variational energies compared to prominent SCI methods such as SCHI [10] and ASCI [12]. After applying an orthonormal rotation to the orbitals—an operation that leaves the exact ground-state energy invariant—the NNBF energy further improves and is in excellent agreement with FCIQMC [38] as well as the perturbatively corrected values from SCHI and ASCI. This performance underscores the NNBF’s capacity for accurately modeling strongly correlated systems. In contrast, other NQS methods have reported energy differences of several mHa for this system [35], further emphasizing the robustness of the NNBF ansatz and the efficacy of our algorithmic enhancements.
Moreover, we performed all-electron calculations for the N2 molecule at selected bond lengths using the cc-pVDZ basis set. The data in Table 3 demonstrate that NNBF not only surpasses the accuracy of conventional quantum chemistry methods up to CCSDTQ but also competes effectively with state-of-the-art Density Matrix Renormalization Group (DMRG) calculations using a bond-dimension of $m=2000$ [39]. These findings indicate that NNBF is among the first NQS approaches capable of tackling such complex, strongly correlated chemical systems with high accuracy. This success opens new avenues for optimizing large-scale NQS wavefunctions more efficiently and effectively in challenging chemical systems.
Method (Parameter) Variational energy Total energy
HF -108.954125 -
CISD -109.242435 -
CCSD -109.263394 -
CCSD(T) - -109.275256
SHCI-HF (Ndets=37593)[10] -109.2692 -109.2769
ASCI-HF (Ndets=10000)[12] -109.26419 -109.27687
ASCI-HF (Ndets=30000)[12] -109.26936 -109.27691
ASCI-HF (Ndets=100000)[12] -109.27335 -109.27698
ASCI-NOR (Ndets=10000)[12] -109.26837 -109.27708
ASCI-NOR (Ndets=100000)[12] -109.27522 -109.27699
ASCI-NOR (Ndets=300000)[12] -109.27638 -109.27699
FCIQMC - HF[38] - -109.2767(1)
NNBF-HF (|$\absolutevalue{\mathcal{S}}$ |=16384, |$\absolutevalue{\mathcal{V}}$ |=4096) -109.276642(36) -
NNBF-NOCCSD (|$\absolutevalue{\mathcal{S}}$ |=16384, |$\absolutevalue{\mathcal{V}}$ |=4096) -109.276911(54) -